# Anonymization Prompt Learning for Facial Privacy-Preserving Text-to-Image Generation

### v1

@@ 1 Introduction
The introduction of text-to-image diffusion models ldm ; dalle3 ; imagen marks a significant milestone in the field of image generation. These models enable the creation of images known for their high resolution, realism, and close alignment with text descriptions. This progress sets new standards for artificial content, and has led to online text-to-image services dalle3 providing free access to the general public.
@@
However, the ability to forge realistic content with such high degrees of freedom has raised wide social concerns. Among the various problems discussed, the generation of identifiable facial images antidreambooth ; hyperdreambooth ; imma draws particular attention, as such realistic deepfake generation could lead to violation of privacy. Multiple incidents have occurred recently where text-to-image models were used to create fake images of public figures, which were often indistinguishable from real photos without careful scrutiny. It is anticipated that future iterations of text-to-image models will only intensify the issue, empowering users to realistically create any individual within any setting solely based on textual descriptions.
@@
In wake of this problem, online service providers develop content moderation systems to avoid undesired face generation. For example, OpenAI’s online image generation system DALLE-3 dalle3 employs a large language model to reject the generation of specific individuals. While effective, such measures introduce extra steps in processing, raising the cost of maintaining a service. To avoid such complicated moderation schemes, researchers also probe into the possibility of removing undesired content from generative models with concept editing forgetmenot ; erase ; ablating . These methods aim at forcing the model to forget sensitive concepts like personal identities forgetmenot ; erase . However, most existing methods are constrained to removing only a handful of concepts without severely damaging generation quality. This inability to process more concepts is potentially due to their modification of all or a large subset of network parameters. Consequently, they are insufficient for defending against deepfake image generation, which involves an unlimited number of facial identities.
@@
In this paper, we propose Anonymization Prompt Learning (APL), a simple approach that addresses the deepfake generation issue. Unlike existing methods, APL achieves this without relying on additional content moderation modules or extensive editing of model parameters. Specifically, we introduce a learnable soft prompt prefix that is prepended to any prompt input for text-to-image models. Given an identity-specific prompt, we train the prefix such that it forces the diffusion model into generating facial images with inaccurate identities, as demonstrated in fig. 1. This is achieved by denoising facial images so that the output aligns less with the requested identity, but closely with a prompt that describes the person’s basic attribute information. We further train the prefix to be inactive for non-identity-specific prompts by aligning the outputs with and without the prefix, thereby maintaining the model’s generation quality. Collectively, a text-to-image diffusion model equipped with Anonymization Prompt (AP) functions almost identically to its original version, except for autonomously generating anonymized images when prompted to depict specific individuals.
@@
Figure 1: Images generated by Realistic Vision v5.1 realisticvision when prompted with names of public figures. The base model generates highly convincing photos of given identities (top row), which may lead to malicious use. When inserting our Anonymization Prompt (AP) to the model (bottom row), the accuracies of generated identities are significantly reduced.
@@
We quantitatively evaluate the anonymization performance with a face recognition model, which measures the distance between generated identities and the ground truths. Results show that our method facilitates the generation of identities significantly less similar to the ground truth identity. Moreover, we establish that the any identities can be anonymized, including those not seen during training. In addition, no significant degradation is observed on generation quality and text fidelity according to commonly-used metrics for image generation evaluation.
@@
We further reveal several advantages of learning a prompt as opposed to training the entire model. Firstly, an Anonymization Prompt learned on one text-to-image diffusion model can be straightforwardly transferred to another model, and demonstrate competitive anonymization performance. This entails a plug-and-play property of our method, facilitating its applicability across a wide array of platforms and scenarios. Secondly, AP remains effective after model fine-tuning, as it can always be integrated to the model following the training process and continue to function.
@@
(1) We propose Anonymization Prompt Learning (APL), a novel approach to mitigate the issue of deepfake generation by text-to-image models. This is achieved through a learnable soft prompt prefix that leads to the generation of anonymized images even for prompts of specific identities.
(2) We ensure the Anonymization Prompt does not affect with the model’s basic generative quality, so that online service providers can integrate the approach into their existing platforms without worrying about performance degradation.
(3) Extensive experiments demonstrate the efficacy of the Anonymization Prompt (AP) in producing anonymized images when models are prompted for specific individuals, without degrading image quality or text fidelity for non-identity-specific requests. Furthermore, the AP exhibits a plug-and-play property that allows it to be easily transferred and integrated into various text-to-image models.
@@ 2.1 Text-to-Image Diffusion Models
The field of image generation has taken a leap forward thanks to the introduction of diffusion models ddpm ; ddim ; adm . Through iterative denoising steps, these models implement a mapping from random noise to the real image distribution, achieving superior performances in terms of image quality and diversity. More recently, Latent Diffusion Model (LDM)  ldm is proposed to operate the diffusion process in a compressed latent space. This practice reduces computational costs and enables the introduction of conditional inputs, such as text descriptions. Taking inspirations from LDM, current text-to-image diffusion models are celebrated for their ability to produce images of high resolution, realism, and close adherence to text descriptions. These advancements have resulted in huge successes like Stable Diffusion ldm ; sdxl , Midjourney midjourney , and DALLE-3 dalle3 , which offer public access to either model parameters or through open API services. Such capabilities enable users to produce an extensive range of realistic creations tailored to their specific requests.
While diffusion models enable numerous creative applications, their utilization raises wide societal concerns regarding potential malicious uses. Particularly, text-to-image models faithfully following textual inputs can be used for inappropriate generations. Among many concerns, one major issue is the generation of high-quality, identifiable facial images, which leads to concerning privacy issues. Several work antidreambooth ; hyperdreambooth ; imma confirm the capability of models to generate accurate identities when prompted with specific individuals. We target the face generation capability of these general-purposed text-to-image models, aiming to moderate them towards failure at identity-specific generation, but preserving abilities to generate high-quality non-identity-specific content.
@@ 2.2 Content Moderation for Text-to-Image Generation
Online text-to-image service providers employ additional pre-processing or post-processing modules to reject generations deemed inappropriate, including images depicting specific individuals dalle3 . Meanwhile, researchers of concept editing methods erase ; forgetmenot ; ablating ; uce explore the modification of text-to-image models themselves to prevent the generation of designated concepts. ESD erase leverages a frozen teacher model to provide unconditional guidance for diffusion models towards forgetting of concepts. Concept Ablation ablating proposes model-based and noise-based ablation that both guide the predictions of diffusion models towards pre-defined anchor concepts. Forget-me-not forgetmenot modifies the cross-attention layers of diffusion models to minimize the attention scores between designated concepts and accurate generations. These methods primarily target a limited number of concepts to remove, and tend to suffer from catastrophic forgetting uce when the number of target concepts increases. This paper demonstrates that a learnable prompt prefix allows for the anonymization of an unlimited number of facial identities, without significantly impacting the quality of generations.
@@ 2.3 Face Anonymization
Face anonymization blur ; pixel ; deepprivacy ; ciagan ; ldfa ; attranon strives to edit existing facial images to remove identity information, thereby protecting facial privacy. Earlier methods blur ; pixel use blurring or pixelation to completely obscure faces. Recent approaches deepprivacy ; ciagan ; ldfa ; attranon utilize generative models to replace existing faces with others that have different identities. Deepprivacy deepprivacy and CIAGAN ciagan use Generative Adversarial Networks gan conditioned on face landmarks to guide accurate anonymizations. LDFA ldfa leverages the powerful inpainting capability of diffusion models to enhance face anonymization for self-driving applications. A recent work attranon proposes to preserve facial attribute information in anonymization with the guidance of a face attribute detection model.
@@ 3.1 Overview
We propose Anonymization Prompt Learning (APL), a simple procedure prevents the generation of identity-specific images in text-to-image diffusion models. We train an Anonymization Prompt (AP) that, by default, is attached to any prompt inputs to diffusion models. The AP is inactivate for most prompt inputs, but forces the diffusion model to generate anonymized images when prepended to prompts specifying facial identities.
In this section, we introduce the idea of learning to anonymize for diffusion models, and provide information on how to preserve prompt fidelity and generation quality of diffusion models with AP introduced. A detailed pipeline of APL is presented in fig. 2.
Figure 2: Pipeline of Anonymization Prompt Learning (APL). At any timestep, the diffusion model denoises images separately with two different prompts. For ID-specific inputs, the Anonymization Prompt (AP) is prepended to prompts of names, and is trained to produce similar noise predictions to that of corresponding attribute descriptions. Similarly, For non-ID-specific inputs, the AP is trained to be ineffective, avoiding generation quality degradation. Both the denoising U-Net and the text encoder are fixed during training.
@@ 3.2 Text-to-Image Diffusion Models
Diffusion models ddpm implement a mapping from random noise to the image distribution by performing a multi-step image denoising. This is achieved by training the model to perform image denoising at different time steps with different noise levels:
@@
where $\epsilon_{t}$ is a sampled Gaussian noise and $t$ is the denoising time step. In the context of Latent Diffusion Models ldm , $x_{t}$ is obtained by compressing real images into a latent space and subsequently adding noises. For text-to-image models, $c$ denotes the textual description guiding the image generation, and $\tau$ is implemented by a text encoder.
@@ 3.3 Prompt Learning for Diffusion Models
where $\oplus$ denotes a concatenation of the AP and the tokenized text. As we will demonstrate, multiple advantages emerge in performing prompt learning instead of training entire models, including a plug-and-play property that enables transferring AP between different text-to-image models.
@@ 3.4 Learning to Anonymize
In existing moderation methods for diffusion models erase ; ablating ; uce , knowledge concerning target concepts are completely removed. However, for the case of face anonymization in text-to-image models, we endeavor to strike a balance between producing completely irrelevant images and retaining original facial attributes, following recent works in face anonymization attranon . In other words, when users request the images of a specific individual, we expect the model to return natural facial images with inaccurate identities, but share similar attributes to the requested identity. This can be especially valuable, as completely altering a person’s appearance could introduce bias, misrepresentation, and degraded user experience.
@@
For example, for $c_{1}$ = "Albert Einstein", we have $c_{2}$ = "A male German-born physicist, bushy mustache, unkempt hair, deep-set eyes. " We note that prompting the model with $c_{2}$ alone almost never leads to the identity specified by $c_{1}$. Therefore, guiding the model to generate images of $c_{2}$ removes the identity, but lead to someone with similar facial attributes that fulfills the description of $c_{2}$. Concretely, we aim at aligning the conditional probabilities $P(x|c_{1})$ and $P(x|c_{2},\rm{not}\ c_{1})$. Following existing work in compositional generalization compositional_energy ; compositional_diffusion , the latter can be approximated with
$P(x|c_{2},\mathrm{not}\ c_{1})\propto P(x|c_{2})(\frac{P(x|c_{2})}{P(x|c_{1})})^{\alpha}.$ (4)
where $\alpha$ is a weight parameter that controls the strength of anonymization. We may subsequently derive the target score function
$\hat{\epsilon}_{\theta}(x,\tau(c_{1}),t)=\epsilon_{\theta}(x,\tau(c_{2}),t)+\alpha(\epsilon_{\theta}(x,\tau(c_{2}),t)-\epsilon_{\theta}(x,\tau(c_{1}),t)).$ (5)
We align score functions modified by Anonymization Prompt with the above target score function:
$L_{ID}=E_{(x,c_{1},c_{2})\sim S_{ID},t}[||\epsilon_{\theta}(x,\tau_{AP}(c_{1}),t)-\hat{\epsilon}_{\theta}(x,\tau(c_{1}),t)||_{2}^{2}].$ (6)
Note that only the first term with Anonymization Prompt backpropagates gradients. Intuitively, $L_{ID}$ guides identifiable images produced by $c_{1}$ to move along the direction between $c_{1}$ and $c_{2}$ to align with the semantics of $c_{2}$, a non-identity-specific prompt. We empirically find this direction of anonymization to generalize very well, as models are able to generate anonymized identities for identities unseen during training.
@@ 3.5 Quality Preservation
In practice, we expect text-to-image service providers to employ Anonymization Prompt by prepending it to any prompts that enter the model. AP then forces the model to automatically produce anonymized identities, therefore saving the demand of additional content moderation modules. To ensure the practicality of our method, it is critical that it does not degrade the base model’s performance for non-identity-specific prompts, including those describing unspecified identities ("A person") or non-facial subjects ("An apple"). The aim is for APL to perform anonymization only when necessary, without compromising the model’s ability to generate high-quality images across various prompts.
We reveal that given the formulation of $L_{ID}$ above, it is very simple to regularize Anonymization Prompt Learning for above purposes. Specifically, we adopt the LAION-2B laion dataset to curate a regularization dataset $S_{Reg}=\{(x^{i},c_{1}^{i},c_{2}^{i})\}_{i=1}^{N}$ , which shares the data structure as $S_{ID}$. For each text-image pair $(c,x)$ in LAION, we curate a sample for $S_{Reg}$ by setting $x^{i}=x$ and $c_{1}=c_{2}=c$. We train on these samples with a similar loss function introduced above:
@@
Effectively, we train the Anonymization Prompt to be aware of the prompts it is concatenated to. When prompts involve facial identities, AP enforces face anonymization. Otherwise, AP is trained to be ineffective, as we guide the model to align the noise predictions for $\tau_{AP}(c)$ and $\tau(c)$.

### v2

@@ Abstract
keywords:
Facial Privacy, Text-to-Image Generation, Prompt Learning, Deepfakes
@@ 1 Introduction
The introduction of text-to-image diffusion models [1, 2, 3] marks a significant milestone in the field of image generation. These models enable the creation of images known for their high resolution, realism, and close alignment with text descriptions. This progress sets new standards for artificial content, and has led to online text-to-image services [2] providing free access to the general public.
@@
However, the ability to forge realistic content with such high degrees of freedom has raised wide social concerns. Among the various problems discussed, the generation of identifiable facial images [4, 5, 6] draws particular attention, as such realistic deepfake generation could lead to violation of privacy. Multiple incidents have occurred recently where text-to-image models were used to create fake images of public figures, which were often indistinguishable from real photos without careful scrutiny. It is anticipated that future iterations of text-to-image models will only intensify the issue, empowering users to realistically create any individual in any setting solely based on textual descriptions.
@@
In wake of this problem, online service providers develop content moderation systems to avoid undesired face generation. For example, OpenAI’s online image generation system DALLE-3 [2] employs a large language model to reject the generation of specific individuals. While effective, such measures introduce extra steps in processing, raising the cost of maintaining a service. To avoid such complicated moderation schemes, researchers also probe into the possibility of removing undesired content from generative models with concept editing [7, 8, 9]. These methods aim at forcing the model to forget sensitive concepts like personal identities [7, 8]. However, most existing methods are constrained to removing only a handful of concepts without severely damaging generation quality. This inability to process more concepts is potentially due to their modification of all or a large subset of network parameters. Consequently, they are insufficient for defending against deepfake image generation, which involves an unlimited number of facial identities.
@@
In this paper, we propose Anonymization Prompt Learning (APL), a simple approach that addresses the deepfake generation issue. Unlike existing methods, APL achieves this without relying on additional content moderation modules or extensive editing of model parameters. To accomplish this, we introduce a learnable soft prompt prefix that is prepended to any prompt input for text-to-image models. Given an identity-specific prompt, we train the prefix such that it forces the diffusion model into generating facial images with inaccurate identities, as demonstrated in fig. 1. This is achieved by denoising facial images so that the output aligns less with the requested identity, but closely with a prompt that describes the person’s prominent facial attribute information. We further train the prefix to be inactive for prompts that do not specify certain identities by aligning the outputs with and without the prefix, thereby maintaining the model’s generation quality. Collectively, a text-to-image diffusion model equipped with the Anonymization Prompt functions almost identically to its original version, except for autonomously generating anonymized images when prompted to depict specific individuals.
@@
Figure 1: Images generated by Realistic Vision v5.1 [10] when prompted with names of public figures. The base model generates highly convincing photos of given identities (top row), which may lead to malicious use. When inserting our Anonymization Prompt (AP) to the model (bottom row), the accuracies of generated identities are significantly reduced.
@@
We quantitatively evaluate the anonymization performance with a face recognition model, which measures the distance between generated identities and the ground truths. Results show that our method facilitates the generation of identities significantly less similar to the ground truth identity. We establish that any identities can be anonymized, including those not seen during training. In addition, no significant degradation is observed on generation quality and text fidelity according to commonly-used metrics for evaluating image generation models.
@@
We further reveal several advantages of learning a prompt as opposed to training the entire model. Firstly, an Anonymization Prompt learned on one text-to-image diffusion model can be straightforwardly transferred to another model, and demonstrate competitive anonymization performance. This entails a plug-and-play property of our method, facilitating its applicability across a wide array of platforms and scenarios. Secondly, the prompt remains effective after model fine-tuning, as it can always be integrated to the model following the training process and continue to function.
@@
(1) We propose Anonymization Prompt Learning (APL), a novel approach to mitigate the issue of deepfake generation by text-to-image models. This is achieved through a learnable soft prompt prefix that leads to the generation of anonymized images even for prompts specifying certain identities.
(2) We ensure the Anonymization Prompt does not affect the model’s basic generative quality, so that online service providers can integrate the approach into their existing platforms without worrying about performance degradation.
(3) Extensive experiments demonstrate the efficacy of the Anonymization Prompt in producing anonymized images when models are prompted for specific individuals, without degrading image quality or text fidelity for non-identity-specific requests. Furthermore, the Anonymization Prompt exhibits a plug-and-play property that allows it to be easily transferred and integrated into various text-to-image models.
@@ 2.1 Text-to-Image Diffusion Models
The field of image generation has taken a leap forward thanks to the introduction of diffusion models [11, 12, 13]. Through iterative denoising steps, these models implement a mapping from random noise to the real image distribution, achieving superior performances in terms of image quality and diversity. More recently, Latent Diffusion Model (LDM)  [1] is proposed to operate the diffusion process in a compressed latent space. This practice reduces computational costs and enables the introduction of conditional inputs, such as text descriptions. Taking inspirations from LDM, current text-to-image diffusion models are celebrated for their ability to produce images of high resolution, realism, and close adherence to text descriptions. These advancements have resulted in huge successes like Stable Diffusion [1, 14], Midjourney [15], and DALLE-3 [2], which offer public access to either model parameters or through open API services. Such capabilities enable users to produce an extensive range of realistic creations tailored to their specific requests.
While diffusion models enable numerous creative applications, their utilization raises wide societal concerns regarding potential malicious uses. Particularly, text-to-image models faithfully following textual inputs can be used for inappropriate generations. Among many concerns, one major issue is the generation of high-quality, identifiable facial images, which leads to concerning privacy issues. Several work [4, 5, 6] confirm the capability of models to generate accurate identities when prompted with specific individuals. We target the face generation capability of these general-purposed text-to-image models, aiming to moderate them towards failure at identity-specific generation, but preserving abilities to generate high-quality non-identity-specific content.
@@ 2.2 Content Moderation for Text-to-Image Generation
Online text-to-image service providers employ additional pre-processing or post-processing modules to reject generations deemed inappropriate, including images depicting specific individuals [2]. Meanwhile, researchers of concept editing methods [8, 7, 9, 16] explore the modification of text-to-image models themselves to prevent the generation of designated concepts. ESD [8] leverages a frozen teacher model to provide unconditional guidance for diffusion models towards forgetting of concepts. Concept Ablation [9] proposes model-based and noise-based ablation that both guide the predictions of diffusion models towards pre-defined anchor concepts. Forget-me-not [7] modifies the cross-attention layers of diffusion models to minimize the attention scores between designated concepts and accurate generations. These methods primarily target a limited number of concepts to remove, and tend to suffer from catastrophic forgetting [16] when the number of target concepts increases. This paper demonstrates that a learnable prompt prefix allows for the anonymization of an unlimited number of facial identities, without significantly impacting the quality of generations.
@@ 2.3 Face Anonymization
Face anonymization [17, 18, 19, 20, 21, 22] strives to edit existing facial images to remove identity information, thereby protecting facial privacy. Earlier methods [17, 18] use blurring or pixelation to completely obscure faces. Recent approaches [19, 20, 21, 22] utilize generative models to replace existing faces with others that have different identities. Deepprivacy [19] and CIAGAN [20] use Generative Adversarial Networks [23] conditioned on face landmarks to guide accurate anonymizations. LDFA [21] leverages the powerful inpainting capability of diffusion models to enhance face anonymization for self-driving applications. A recent work [22] proposes to preserve facial attribute information in anonymization with the guidance of a face attribute detection model.
@@ 3.1 Overview
We propose Anonymization Prompt Learning (APL), a simple procedure that prevents the generation of identity-specific images in text-to-image diffusion models. We train an Anonymization Prompt that, by default, is attached to any prompt inputs to diffusion models. The Anonymization Prompt is inactivate for most prompt inputs, but forces the diffusion model to generate anonymized images when prepended to prompts specifying facial identities.
In this section, we introduce the idea of learning to anonymize for diffusion models, and provide information on how to preserve prompt fidelity and generation quality of diffusion models with APL introduced. A detailed pipeline of APL is presented in fig. 2.
Figure 2: Pipeline of Anonymization Prompt Learning (APL). At any timestep, the diffusion model denoises images separately with two different prompts. For ID-specific inputs from $S_{ID}$, the Anonymization Prompt (AP) is prepended to prompts of names, and is trained to produce similar noise predictions to that of corresponding attribute descriptions. Similarly, For non-ID-specific inputs from $S_{Reg}$, the Anonymization Prompt is trained to be ineffective, avoiding generation quality degradation. Both the denoising U-Net and the text encoder are fixed during training.
@@ 3.2 Text-to-Image Diffusion Models
Diffusion models [11] implement a mapping from random noise to the image distribution by performing a multi-step image denoising. This is achieved by training the model to perform image denoising at different time steps with different noise levels:
@@
where $\epsilon_{t}$ is a sampled Gaussian noise and $t$ is the denoising time step. In the context of Latent Diffusion Models [1], $x_{t}$ is obtained by compressing real images into a latent space and subsequently adding noises. For text-to-image models, $c$ denotes the textual description guiding the image generation, and $\tau$ is implemented by a text encoder.
@@ 3.3 Prompt Learning for Diffusion Models
where $\oplus$ denotes a concatenation of the prompt and the tokenized text. As we will demonstrate, multiple advantages emerge in performing prompt learning instead of training entire models, including a plug-and-play property that enables transferring the prompt between different text-to-image models.
@@ 3.4 Learning to Anonymize
In existing moderation methods for diffusion models [8, 9, 16], knowledge concerning target concepts are completely removed. However, for the case of face anonymization in text-to-image models, we endeavor to strike a balance between producing completely irrelevant images and retaining original facial attributes, following recent works in face anonymization [22]. In other words, when users request the images of a specific individual, we expect the model to return natural facial images with inaccurate identities, but share similar attributes to the requested identity. This can be especially valuable, as completely altering a person’s appearance could introduce bias, misrepresentation, and degraded user experience.
@@
For example, for $c_{1}$ = "Albert Einstein", we have $c_{2}$ = "A male German-born physicist, bushy mustache, unkempt hair, deep-set eyes. " We note that prompting the model with $c_{2}$ alone almost never leads to the identity specified by $c_{1}$. Therefore, guiding the model to generate images of $c_{2}$ removes the identity, but lead to someone with similar facial attributes that fulfills the description of $c_{2}$. Concretely, we aim at aligning the conditional probabilities $P(x|c_{1})$ and $P(x|c_{2},\rm{not}\ c_{1})$. Following existing work in compositional generalization [24, 25], the latter can be approximated with
$P(x|c_{2},\mathrm{not}\ c_{1})\propto P(x)\frac{P(x|c_{2})}{P(x|c_{1})}.$ (4)
We compose the score function in such a way that approximates above probability:
$\hat{\epsilon}_{\theta}(x,\tau(c_{1}),t)=\epsilon_{\theta}(x,\tau(c_{2}),t)+\alpha(\epsilon_{\theta}(x,\tau(c_{2}),t)-\epsilon_{\theta}(x,\tau(c_{1}),t)),$ (5)
where $\alpha$ is a weight parameter that controls the strength of the anonymization.
Finally, to enforce the maximization of the probability, we align the score functions modified by the Anonymization Prompt with this target score function:
$L_{ID}=E_{(x,c_{1},c_{2})\sim S_{ID},t}[||\epsilon_{\theta}(x,\tau_{AP}(c_{1}),t)-\hat{\epsilon}_{\theta}(x,\tau(c_{1}),t)||_{2}^{2}],$ (6)
Note that the gradients only flow through the first term with the Anonymization Prompt. Intuitively, $L_{ID}$ guides identifiable images produced by $c_{1}$ to move along the direction between $c_{1}$ and $c_{2}$ to align with the semantics of $c_{2}$, a non-identity-specific prompt. We empirically find this direction of anonymization to generalize very well, as models are able to generate anonymized identities for identities unseen during training.
@@ 3.5 Quality Preservation
In practice, we expect text-to-image service providers to employ the Anonymization Prompt by prepending it to any prompts that enter the model. This way, no extra module is needed for examining the content of prompts. For identity-specific prompts, the Anonymization Prompt forces the model to automatically produce anonymized identities. To ensure practicality, it is critical that it does not degrade the base model’s performance for non-identity-specific prompts, including those describing unspecified identities ("A person") or non-facial subjects ("An apple"). The aim is for APL to perform anonymization only when necessary, without compromising the model’s ability to generate high-quality images across various prompts.
We reveal that given the formulation of $L_{ID}$ above, it is very simple to regularize Anonymization Prompt Learning for above purposes. Specifically, we adopt the LAION-2B [26] dataset to curate a regularization dataset $S_{Reg}=\{(x^{i},c_{1}^{i},c_{2}^{i})\}_{i=1}^{N}$ , which shares the data structure as $S_{ID}$. For each text-image pair $(c,x)$ in LAION, we curate a sample for $S_{Reg}$ by setting $x^{i}=x$ and $c_{1}=c_{2}=c$. We train on these samples with a similar loss function introduced above:
@@
Effectively, we train the Anonymization Prompt to be aware of the prompts it is concatenated to. When prompts involve facial identities, the Anonymization Prompt enforces face anonymization. Otherwise, it is trained to be ineffective, as we guide the model to align the noise predictions for $\tau_{AP}(c)$ and $\tau(c)$.